OpenStudy (anonymous):

i need help -12(3+14)=(-12+3)+14

OpenStudy (anonymous):

What is the question asking? There is no variable, so i am assuming it is asking whether or not these are equal. -12(3+14)=(-12+3)+14 -12(17)=(-9)+14 -204=5 It is NOT equal.

OpenStudy (anonymous):

If it is asking what property is being represented, it would be the associative property, I think.